Tthe perimeter of a parallelogram is 62 inches. If the length of one side is 20 inches, what is the side adjacent to it?

Answer by checkley79(3341)

P=2L+2W
62=2*20+2W
62=40+2W
2W=62-40
2W=22
W=22/2
W=11 ANS.
PROOF:
62=2*20+2*11
62=40+22
62=62